Net Present Value Method and Present Value Index

Diamond and Turf Inc. is considering an investment in one of two machines. The sewing machine will increase productivity from sewing 160 baseballs per hour to sewing 288 per hour. The contribution margin per unit is $0.48 per baseball. Assume that any increased production of baseballs can be sold. The second machine is an automatic packing machine for the golf ball line. The packing machine will reduce packing labor cost. The labor cost saved is equivalent to $28 per hour. The sewing machine will cost $427,700, have a 10-year life, and will operate for 1,400 hours per year. The packing machine will cost $151,900, have a 10-year life, and will operate for 1,200 hours per year. Diamond and Turf seeks a minimum rate of return of 12% on its investments.

Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.353 2.991
6 4.917 4.355 4.111 3.785 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

a. Determine the net present value for the two machines. Use the table of present values of an annuity of $1 above. Round to the nearest dollar.

Sewing Machine Packing Machine
Present value of annual net cash flows $ $
Amount to be invested $ $
Net present value $ $

b. Determine the present value index for the two machines. If required, round your answers to two decimal places.

Sewing Machine Packing Machine
Present value index

c. If Diamond and Turf has sufficient funds for only one of the machines and qualitative factors are equal between the two machines, in which machine should it invest? (If both present value indexes are the same, either machine will grade as correct.)
Packing Machine
Answer #1
Annual cash flows:
Sewing Machine 86016 =(288-160)*0.48*1400
Packing Machine 33600 =28*1200
a
Sewing Machine Packing Machine
Present value of annual net cash flows 485990 189840
Amount to be invested 427700 151900
Net present value 58290 37940
b
Sewing Machine Packing Machine
Present value index 1.14 1.25
=485990/427700 =189840/151900
c
Packing Machine
Workings:
Sewing Machine Packing Machine
Present value of annual net cash flows =86016*5.650 =33600*5.650